OpenStudy (anonymous):

How do I solve radical x^2-9 plus x+1=0

OpenStudy (anonymous):

\[\sqrt{x ^{2}+9} +x +1=0\]

OpenStudy (abb0t):

Now it's plus?

OpenStudy (anonymous):

sorry, I wrote the question wrong

OpenStudy (abb0t):

\(\sqrt{x^2} \sqrt{9} + x +1=0\)

OpenStudy (anonymous):

would it be \[3\sqrt{x ^{2}}=-x-1\]

OpenStudy (phi):

if the problem is \[ \sqrt{x ^{2}+9} +x +1=0 \] this is equivalent (add -(x+1) to both sides) to \[ \sqrt{x ^{2}+9} = -(x +1) \] square both sides \[ x^2+9 = x^2 +2x +1 \\ 8=2x\\x=4\] plug in x=4 into the original equation: \[ \sqrt{16+9} +4 +1=0 \] in these type problems, we take the positive square root so this is \[ 5 + 5=0 \\10=0\] that means there is no solution

OpenStudy (anonymous):

Thank you! @phi